OpenStudy (anonymous):

Simplify the radical expression sqrt{7}+3 sqrt{7} a)2 sqrt{7} b)4 sqrt{14} c)4 sqrt{7} d)2 sqrt{14}

OpenStudy (phi):

if you have 1 square root of 7 and add 3 more square root of 7's how many sqr(7) do you have ?

OpenStudy (anonymous):

if you work it out completly it is 11.37673

OpenStudy (anonymous):

14

OpenStudy (anonymous):

so would it be D?

OpenStudy (phi):

try a simpler problem first: you have 1 dog and add 3 more dogs, how may dogs? you have 1 apple and add 3 more apples, how many apples ? you have 1 sqr(7) and add 3 more sqr(7), how many sqr(7) ?

OpenStudy (anonymous):

Oh, it's 4! Haha oops

OpenStudy (phi):

and you write it as \[ 4 \sqrt{7} \] which means 4 sqr(7) or \[ \sqrt{7}+ \sqrt{7}+ \sqrt{7}+ \sqrt{7} \] either way, 4 of them

OpenStudy (anonymous):

Okay, thanks so much!! Can you help me simply two other expressions?

OpenStudy (anonymous):

lololol

OpenStudy (anonymous):

\[2\sqrt{6+3\sqrt{96}}\]

OpenStudy (anonymous):

Thats the next one I need help with

OpenStudy (anonymous):

@phi

OpenStudy (phi):

you should make it a new post, otherwise these things get a bit long. but the first thing I would do is find all the prime factors of 96 can you do that ?

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

1, 2, 3, 4, 6, 8, 12, 16, 24, 32,48, and 96. Right?

OpenStudy (anonymous):

\[50\sqrt{1}\]

OpenStudy (anonymous):

Are you sure its not \[50\sqrt{6}\]? beccause I don't have the answer choice of \[50\]

OpenStudy (anonymous):

\[50\sqrt{1}\]

OpenStudy (phi):

just the prime factors (prime numbers) that when you multiply them you get 96 you get 2*2*2*2*2*3

OpenStudy (phi):

btw, is the problem \[\sqrt[2]{6+\sqrt[3]{96}}\] ?

OpenStudy (anonymous):

Ohh, okay.

OpenStudy (anonymous):

So then what??

OpenStudy (phi):

in other words is the cube root of 96 or 3 times the square root of 96 ? \[ \sqrt[3]{96} \text{ or } 3 \sqrt{96} ? \]

OpenStudy (anonymous):

No it's not

OpenStudy (phi):

? which is it ? can you post a screen shot ?

OpenStudy (anonymous):

I have it posted up a little higher

OpenStudy (anonymous):

I can't screenshot, but i will post the equation again quick. Hang on.

OpenStudy (anonymous):

\[2\sqrt{6+3\sqrt{96}}\]

OpenStudy (phi):

yes, I see the expression \[ 2\sqrt{6+3\sqrt{96}} \] but I want to know if you really mean 3 times sqrt(96) or cube root of 96 (and the same for the 2)

OpenStudy (anonymous):

No, that's it.

OpenStudy (phi):

ok. the sqr(96) is the same as sqr(2*2 * 2*2 * 2 *3) "pull out pairs" from the square root, and "toss one of each pair" \[ \sqrt{2\cdot 2 \cdot 2\cdot 2\cdot 2 \cdot 3}= 2\cdot 2 \sqrt{2 \cdot 3} \] or \[ 4 \sqrt{6} \]

OpenStudy (phi):

and the problem is now \[ 2\sqrt{6+ 3\cdot 4 \sqrt{6}} \] or \[ 2\sqrt{6+ 12 \sqrt{6}} \]

OpenStudy (anonymous):

you mean \[14\sqrt{6}\]?

OpenStudy (anonymous):

nvm, but I have the answer now

OpenStudy (anonymous):

I need help with one more but I will make a different post of it and tag you in it

OpenStudy (phi):

what is the answer for this one ? (out of curiosity)

OpenStudy (anonymous):

I think it's c. :)